# Probing light sterile neutrinos in left-right symmetric models with displaced vertices and neutrinoless double beta decay

### v1

@@ 8.2 Type-I seesaw benchmark scenarios
In Fig. 11, we show in the $M_{4}$-$M_{W_{R}}$ plane whether the minimal-mixing contributions or RH-current contributions dominate the sterile-neutrino decay for various $m_{1}$ values. For example, we see that for a large $m_{1}=0.03$ eV, the minimal-mixing contributions dominate the decays of the lightest sterile neutrinos for $M_{4}<0.9$ GeV (around the $\rho$-meson threshold) if $M_{W_{R}}\gtrsim 20$ TeV. As shown in Fig. 10, DV-search experiments such as DUNE, SHiP, and MATHUSLA can probe RH gauge-boson effects in this mass range in type-II seesaw scenarios. This means minimal-mixing contributions can be significant for a relatively large $m_{1}$. Hence, we deem it interesting to further analyze the sensitivity reach in type-I seesaw benchmark scenarios, which we now do.
@@
To maximize the impact of minimal-mixing effects, we set $m_{1}=0.03$ eV. In Fig. 12, we show isocurves where $N_{N}^{\text{obs}}=3$ for the (future) DV-search experiments in the considered type-I seesaw scenario. All the LHC far detectors that probe $\sim 10$ TeV RH gauge-boson masses exhibit similar isocurves compared to the type-II scenario. This is unsurprising as RH contributions dominate at this magnitude of $M_{W_{R}}$. In the case of DUNE, however, the sensitivity in the type-I seesaw scenario is perceptibly higher than in the type-II seesaw scenario. Finally, we note that the results are insensitive to the choice between generalized $\mathcal{P}$-symmetry and $\mathcal{C}$-symmetry scenarios.
@@
Since there are no direct relations between the active- and sterile-neutrino masses in the type-I seesaw scenario, there is less predictive power for the $0\nu\beta\beta$ sensitivity reach compared to type-II seesaw scenarios. Nevertheless, comparing the $0\nu\beta\beta$ sensitivity reach to the far-detector sensitivity reaches is still interesting. We choose the sterile-neutrino masses to be large ($M_{5,6}=100\text{ GeV}\gg M_{B_{s}}$) to be conservative. In Fig. 12, we show the results for current and next-gen experiments in this type-I seesaw scenario. Again the existing $0\nu\beta\beta$ constraints are stringent, but DUNE and SHiP will be more sensitive in parts of the parameter space.
@@
Because of the relatively large $m_{1}$ at 0.03 eV, the minimal-mixing contributions are sufficiently large to be detected in next-generation $0\nu\beta\beta$ experiments for $M_{4}>0.5$ GeV. This is true irrespective of the mass of the RH gauge boson and therefore the green dashed line becomes vertical. However, For $M_{4}<0.5$ GeV, the active-neutrino contributions to the ${}^{136}\text{Xe}$ decay rate can cancel the sterile-neutrino contributions, analogous to the phase cancellations occurring in parts of the gray regions in Fig 8. In this case, the next-generation $0\nu\beta\beta$ experiments will be sensitive to $M_{W_{R}}=25$ TeV roughly.
@@
Figure 12: Predicted sensitivity reach of the next-gen $0\nu\beta\beta$ experiments and the DV searches in the $M_{4}-M_{W_{R}}$ plane in the type-I seesaw scenario setting $m_{1}=0.03$ eV. For the $W_{L}$-$W_{R}$ mixing, we have set $\xi=0.3$ ($\xi=0.0$) in the left (right) panel. We also show existing bounds from the $0\nu\beta\beta$ experiment KamLAND-Zen KamLAND-Zen:2022tow . We show recast bounds obtained from NA62 NA62:2020mcv , T2K T2K:2019jwa , and BEBC WA66:1985mfx ; Barouki:2022bkt .
@@ 9 Conclusions
By performing Monte-Carlo simulations, we have investigated the sensitivity reaches of the following (future) DV experiments: ANUBIS, CODEX-b, DUNE, FACET, FASER(2), MATHUSLA, MoEDAL-MAPP1(2), and SHiP. We have considered $B$-meson, $D$-mesons, and kaon decays, where we have probed scenarios with and without the $W_{L}$-$W_{R}$ mixing. We highlight the sensitivity reach of the DUNE and SHiP experiments, which could probe sterile-neutrino signal events for RH gauge-boson masses $\sim 20$ TeV.

### v2

@@ 8.2 Type-I seesaw benchmark scenarios
In Fig. 11, we show in the $M_{4}$-$M_{W_{R}}$ plane whether the minimal-mixing contributions or RH-current contributions dominate the sterile-neutrino decay for various $m_{1}$ values. For example, we see that for a large $m_{1}=0.03$ eV, the minimal-mixing contributions dominate the decays of the lightest sterile neutrinos for $M_{4}<0.9$ GeV (around the $\rho$-meson threshold) if $M_{W_{R}}\gtrsim 20$ TeV. As shown in Fig. 10, DV-search experiments such as DUNE, SHiP, and MATHUSLA can probe RH gauge-boson effects in this mass range in type-II seesaw scenarios. This means minimal-mixing contributions can be significant for a relatively large $m_{1}$. Hence, we deem it interesting to analyze further the sensitivity reach in type-I seesaw benchmark scenarios, which we now do.
@@
To maximize the impact of minimal-mixing effects, we set $m_{1}=0.03$ eV. In Fig. 12, we show isocurves where $N_{N}^{\text{obs}}=3$ for the (future) DV-search experiments in the considered type-I seesaw scenario. All the LHC far detectors that probe $\sim 10$ TeV RH gauge-boson masses exhibit similar isocurves compared to the type-II scenario. This is unsurprising as RH contributions dominate at this magnitude of $M_{W_{R}}$. However, in the case of DUNE, the sensitivity in the type-I seesaw scenario is perceptibly higher than in the type-II seesaw scenario. Finally, we note that the results are insensitive to the choice between generalized $\mathcal{P}$-symmetry and $\mathcal{C}$-symmetry scenarios.
@@
Since there are no direct relations between the active- and sterile-neutrino masses in the type-I seesaw scenario, there is less predictive power for the $0\nu\beta\beta$ sensitivity reach compared to type-II seesaw scenarios. Nevertheless, comparing the $0\nu\beta\beta$ sensitivity reach to the far-detector sensitivity reaches is still interesting. We choose the sterile-neutrino masses to be large ($M_{5,6}=100\text{ GeV}\gg M_{B_{s}}$) to be conservative. In Fig. 12, we show the results for current and next-gen experiments in this type-I seesaw scenario. Again, the existing $0\nu\beta\beta$ constraints are stringent, but DUNE and SHiP will be more sensitive in parts of the parameter space.
@@
Because of the relatively large $m_{1}$ at 0.03 eV, the minimal-mixing contributions are sufficiently large to be detected in next-generation $0\nu\beta\beta$ experiments for $M_{4}>0.5$ GeV. This is true irrespective of the mass of the RH gauge boson, and therefore, the green dashed line becomes vertical. However, For $M_{4}<0.5$ GeV, the active-neutrino contributions to the ${}^{136}\text{Xe}$ decay rate can cancel the sterile-neutrino contributions, analogous to the phase cancellations occurring in parts of the gray regions in Fig 8. In this case, the next-generation $0\nu\beta\beta$ experiments will be sensitive to $M_{W_{R}}=25$ TeV roughly.
@@
Figure 12: Predicted sensitivity reach of the next-gen $0\nu\beta\beta$ experiments and the DV searches in the $M_{4}-M_{W_{R}}$ plane in the type-I seesaw scenario setting $m_{1}=0.03$ eV. For the $W_{L}$-$W_{R}$ mixing, we have set $\xi=0.3$ ($\xi=0.0$) in the left (right) panel. We also show existing bounds from the $0\nu\beta\beta$ experiment KamLAND-Zen KamLAND-Zen:2022tow . We show recast bounds obtained from PIENU PIENU:2019usb , NA62 NA62:2020mcv , T2K T2K:2019jwa , and BEBC WA66:1985mfx ; Barouki:2022bkt .
@@ 9 Conclusions
By performing Monte-Carlo simulations, we have investigated the sensitivity reaches of the following (future) DV experiments: ANUBIS, CODEX-b, DUNE, FACET, FASER(2), MATHUSLA, MoEDAL-MAPP1(2), and SHiP. We have considered $B$-meson, $D$-meson, and kaon decays, where we have probed scenarios with and without the $W_{L}$-$W_{R}$ mixing. We highlight the sensitivity reach of the DUNE and SHiP experiments, which could probe sterile-neutrino signal events for RH gauge-boson masses $\sim 20$ TeV.